OpenStudy (anonymous):

How would i convert these rectangular coordinates into polar coordinates? (3,pi/2)

OpenStudy (anonymous):

You certain you want to go that way? Those coordinates seem to be more likely to be polar coordinates as they are.

OpenStudy (anonymous):

Crap, you're right, i missed worded the question. Those are indeed polar coordinates and i need help converting them into rectangular coords.

OpenStudy (anonymous):

 For rectangular we need a x value and a y value. x value is \[r \cos (\theta)\] r is the length in polar form. In this case 3. y value is \[rsin(\theta) \]

OpenStudy (anonymous):

Take them to the north/south pole.... eeeyyyyhh eeeeyyyyhhh

OpenStudy (anonymous):

pi/2 = 90 deg 3cos(90) = 0 3sin(90) = 3 rectangular point in (0,3). And it matches up with the picture.

OpenStudy (anonymous):

so i would need to find the cos of pi/2 then multiply it by 3? then find the sin of pi/2 and multiply by 3?

OpenStudy (anonymous):

Right. But be careful when using your calculator. You can use pi/2 when your calculator is in rad mode, but in deg mode you have to use 90.

OpenStudy (anonymous):

right. i would get the same answer if i manually multiplied 3 by the radical form of cos and sin right?

OpenStudy (kainui):

x=r*cos(theta) y=r*sin(theta) To go the other way, r=sqrt(x^2+y^2) theta = arctan(y/x) Think about why this is true.  Just use SOH CAH TOA and the pythagorean theorem. =)

OpenStudy (anonymous):

ok sweet, i guess i was just over complicating cuz the formulas seem really simple now. We thanks guys i appreciate the help alot

OpenStudy (kainui):

Yeah, learning can be kind of fun and satisfying like that haha.